If a farmer did head count and got 200 livestock (and he keeps chicken n pigs), counted the legs and got 540, how many chicken were there?
Answer by fcabanski(1391)

C is chickens and P is pigs. Each animal has one head so C + P = 200.


Chickens have two legs, so for legs chickens add 2*C (2 times the number of chickens.) Pigs have four legs, so for legs pigs add 4*P. 2C + 4P = 540.


From the first equation solve for C: C = 200-P. Substitute that into the second equation.


2(200-P)+4P = 540


400 -2P + 4P = 540


2P = 140


P = 70.


C = 200-70=130.

Hope the solution helped. Sometimes you need more than a solution. Contact fcabanski@hotmail.com for online, private tutoring, or personalized problem solving (quick for groups of problems.)